The length of a rectangle is three times its width.
If the area of the rectangle is 108cm2, what is its perimeter?

Answer by rfer(16322)

x*3x=108
3x^2=108
x^2=108/3
x^2=36
x=sqrt36
x=6
3x=18
---------
2*6+2*18=48 cm